# OpenFactCheck: A Unified Framework for Factuality Evaluation of LLMs

### v1

@@ Abstract
The increased use of large language models (LLMs) across a variety of real-world applications calls for mechanisms to verify the factual accuracy of their outputs. Difficulties lie in assessing the factuality of free-form responses in open domains. Also, different papers use disparate evaluation benchmarks and measurements, which renders them hard to compare and hampers future progress. To mitigate these issues, we propose OpenFactCheck, a unified factuality evaluation framework for LLMs. OpenFactCheck consists of three modules: (i) CustChecker allows users to easily customize an automatic fact-checker and verify the factual correctness of documents and claims, (ii) LLMEval, a unified evaluation framework assesses LLM’s factuality ability from various perspectives fairly, and (iii) CheckerEval is an extensible solution for gauging the reliability of automatic fact-checkers’ verification results using human-annotated datasets. OpenFactCheck is publicly released at https://github.com/yuxiaw/OpenFactCheck.
@@ 1 Introduction
Large language models (LLMs) have demonstrated impressive capabilities in generating naturally-sounding answers over a broad range of human inquiries. However, GPT-4 (OpenAI, 2023) and other text generation models still frequently produce content that deviates from real-world facts (Bang et al., 2023; Borji, 2023; Guiven, 2023). This degrades the system performance and undermines its reliability, representing a significant bottleneck in the deployment (Chuang et al., 2023; Geng et al., 2023).
@@
Many studies have explored evaluating and improving the factuality of LLMs Lee et al. (2022); Chuang et al. (2023); Shi et al. (2023); Chen et al. (2023). Two challenges have been identified for evaluation: (1) it is difficult to assess the factuality of open-domain free-form responses, and (2) different papers use different evaluation datasets and measurements, rendering them hard to compare and hampering future progress Wang et al. (2024). To mitigate these issues, we introduce an open-sourced fact-checking system OpenFactCheck.
@@
Figure 1: Overview of the OpenFactCheck demo system for LLM factuality evaluation with three modules. Green CustChecker: a customized fact-checker to identify factual errors given the outputs of LLMs. Orange LLMEval: a unified LLM factuality evaluator to assess the LLM factual ability from different aspects and then to produce a report to illustrate the weakness and to offer improvement advice. Purple CheckerEval: a fact-checker evaluator and leaderboard to encourage the development of advanced checkers in terms of performance, time and costs.
@@
It includes three modules as shown in Figure 1. CustChecker allows users to customize an automatic fact-checker and to verify free-form documents to alleviate the first problem. A unified LLM factuality evaluation module LLMEval applies seven factuality-specific benchmarks to assess the LLM factuality ability from different aspects and then produces a report to illustrate the weakness and offer improvement advice, tackling the second challenge. We further incorporate CheckerEval that assesses the verification accuracy of fact-checkers, equipped with a leaderboard in terms of accuracy, latency, and costs, aiming to encourage the development of advanced automatic fact-checking systems.
@@
The three modules collaborate and help each other. The results of human verification derived from LLMEval can be used as the benchmark for evaluating the accuracy of automated fact-checkers. Simultaneously, the most effective checker identified in CheckerEval can be deployed for automated fact-checking tasks. Each fact-checker in CheckerEval can be an implementation in CustChecker. Complex user inquiries may be considered as potential candidates included the factuality assessment dataset utilized in LLMEval.
@@
Users can tailor their checkers according to their specific needs, such as domain specialization, cost-effectiveness, or rapid processing, and identify factual errors for both human-written text (a claim or document) and the outputs of LLMs. LLM researchers and practitioners can directly submit their LLM responses to our LLMEval platform by downloading our question set. Subsequently, we conduct evaluations to assess the model’s factual accuracy and to generate a report analyzing the model performance from multiple aspects.
Similarly, developers who seek to evaluate and to compare the efficacy of their fact-checking systems to other ones fairly can upload their checker’s verification outcomes to CheckerEval. Then, our system will show the ranking information in the leaderboard after evaluating under the same measurements.
From the perspective of developers and contributors to this open-source project, we encourage extensive implementation of unique, effective, and robust claim processors, retrievers and verifiers within fact-checking pipelines, collections of challenging questions that LLMs tend to make factual errors, and human-annotated fine-grained verification examples. We believe that this will help to promote and to advance future research on LLM factuality.
To sum, this work investigates three research questions in LLM factuality evaluation:
  • how to effectively identify factual errors in an LLM response;
@@
We initiate an open-source project and develop a preliminary version implementing the three modules, which is anticipated to serve as a stepping stone to facilitate future endeavors in this domain.
@@ 2 Background
Fact-checking is the task of assessing whether claims made in writing are manipulated or true.
@@ 2.1 Fact-checking Systems
Many recent papers have described automatic fact-checking systems used to evaluate the factuality of LLM responses, such as RARR, FactScore, FacTool, CoVe, and Factcheck-GPT Gao et al. (2022); Min et al. (2023); Chern et al. (2023); Dhuliawala et al. (2023); Wang et al. (2023). Each checker has unique characteristics designed for specific domains or scenarios. RARR verifies a document as a whole and can generate an attribution report to explain factual errors. FactScore retrieves evidence from an offline Wikipedia dump mainly for the biography. FacTool is friendly to users with low latency, CoVe completely depends on the capability of LLMs, and Factcheck-GPT has a fine-grained pipeline to localize intermediate errors. Unlike the above work, our aim is to enable the easy customization of a fact-checker according to users’ requirements and application scenarios, e.g., offline settings with a limited budget, by simply clicking dropdowns to choose offline retrievers and verifiers supported by small models without calling APIs.
@@
Despite different designs and implementations of various checkers, they generally consist of three modules: (1) claim processor, which extracts context-independent atomic claims from a document, (2) evidence retriever, which searches related passages from the Internet or database, and then ranks them by relevance, and (3) verifier, which determines the claim/document factuality based on the collected evidence (Guo et al., 2022; Li et al., 2023b; Wang et al., 2024). To this end, we first unify different fact-checking systems into a unified pipeline with the three modules. Then, given a module, users can select a developed module from various implementations or develop one by themselves. In addition, the framework supports easy migration of existing fact-checking systems to our demo.
However, the verification results of automatic fact-checkers are not necessarily accurate. How to evaluate and improve the accuracy of automated fact-checkers is critical. The accuracy of automatic fact-checkers serves as a confidence and a reliability signal for the verification results.
2.2 Automatic Evaluation of Fact-Checking Systems
@@
How accurate are current fact-checking systems? Can they effectively serve as proxies for evaluating the factual accuracy of language models? Existing automatic fact-checking studies often first collect a set of human-annotated (LLM response, extracted claims, factuality of the claims), and then quantify the effectiveness of their systems by comparing the final verification results (i.e., whether a claim or a document is factually true or false) to human-annotated labels Min et al. (2023); Chern et al. (2023); Dhuliawala et al. (2023). Recent work on long-form factuality in LLMs also demonstrates a statistically significant correlation between the outputs by automatic fact-checkers and labels by human annotators (Wei et al., 2024). Thus, we merge four human-annotated LLM factuality datasets including FacTool-QA, FELM-WK, Factcheck-Bench, and HaluEval, and then compare them to the results of automatic fact-checkers to assess the performance of fact-checking systems.
@@ 2.3 LLM Factuality Evaluation
Relevance vs. Factuality There are subtle differences between evaluating LLM’s general performance and factuality. Question answering (QA) datasets over various domains are always used for general performance evaluation (Hendrycks et al., 2021). The focus is on judging whether the question is answered correctly. If the model’s response contains the correct answer, it counts; otherwise, it is void even though the response presents all facts. While research on factuality concentrates more on whether the response presents facts aligning with world knowledge even if some statements were irrelevant to the question.
Therefore, instead of using datasets for general performance assessment, we selected seven datasets that are specifically collected for factuality evaluation. They were selected to cover as diverse potential factual errors as possible, including aspects of vulnerabilities of snowballing hallucinations, awareness of self-uncertainty, robustness to false-premise questions, fresh questions with answers changing fast over time, and free-form responses spanning distinct domains, topics, and tasks.
After evaluations over datasets from different perspectives, an analysis report will be generated to demonstrate the advantage and weaknesses of the model and the potential solutions to improve.
Summary:
Observations above motivate us to integrate these three closely-related components into one platform to facilitate (i) users to flexibly configurate an automatic fact-checking system to verify the factuality of claims and documents, (ii) LLM developers to evaluate LLM’s factuality under the same measurement scale including the task, dataset, and metrics, and (iii) researchers to assess the fact-checkers reliability under the fine-grained annotated benchmarks.
3 Design and Implementation
Our OpenFactCheck platform is implemented by a Python server, a web user interface, and a database, deployed via AWS. The Python backend can also be used as a Python toolkit, allowing easy and flexible development. OpenFactCheck’s design emphasizes two principles: (i) customizability and extensibility for both users and developers, and (ii) compatibility with existing methods and datasets. It consists of three modules: CustChecker, LLMEval, and CheckerEval. Below, we present the detailed design and implementation of each component.
@@ 3.1 CustChecker
CustChecker allows users to customize a fact-checking system by selecting a claim processor, a retriever, and a verifier in web pages. The current version currently supports the following fact-checking systems: RARR, FacTool and Factcheck-GPT (Gao et al., 2022; Chern et al., 2023; Wang et al., 2023). Users input either human-written text or outputs of LLMs into the box (see Figure 4), and then the fact-checker defined above will process and detect factual errors, showing the verification results including evidence, judgment, and explanations.
Figure 2: Pseudo code for the three modules in CustChecker.
@@
Table 1: FactQA: vulnerability that is being evaluated, domain, and size across seven datasets.
@@
We unify diverse fact-checking systems as a procedure with three steps, abstracted into three classes: a claim_processor, a retriever, and a verifier. The instances of the three classes are sequentially linked into a pipeline, solving the following tasks: (1) decomposing a document into atomic claims, (2) collecting relevant passages given a claim, and (3) making a true/false judgment given both the claim and the evidence as input (see Figure 2). We refer to them as task solvers. The implementation of a task solver can be flexible, just ensuring that the input and the output are aligned with the abstract class definitions. For example, evidence can be retrieved by calling SerpAPI or by searching Wikipedia using BM25, but we must return a list of relevant passages given an input claim.
Moreover, task solvers in our pipeline are not hardcoded, but can be configured through a yaml configuration file. Thus, users can combine task-solver implementations from different frameworks (e.g., using Factcheck-GPT’s claim processor, RARR’s retriever, and FacTool’s verifier) and start the verification from any step. For example, users can start from the step of retrieval when the input document does not need decomposition.
This functionality is achieved using a message-passing mechanism that uses a success flag as the message indicates whether the current task solver successfully executes and returns the expected output. The success flag passes through the configured order of the pipeline, guaranteeing that the output of the preceding solver fits the input for the current solver, otherwise error warning will be issued. Practically, the input and the output parameter names for the task solvers are defined in the configuration file. To link different solvers into a pipeline, one only needs to ensure that the current solver output name matches the input name of the succeeding solver. A dictionary format fact-checking-state is kept throughout the pipeline to store all information in the verification.
@@
Inspired by Fairseq (Ott et al., 2019), our framework is designed to be highly extendable by treating any third-party task solvers as plugins. As long as the developed task solvers adhere to our class interface definitions, they can be imported and used in our framework.
In summary, due to the customizable and extendable nature of CustChecker, general users of OpenFactCheck can utilize it as a fully functional application with web-based user interfaces. Advanced developers have the flexibility to use it as a library, allowing them to develop their task solvers or to integrate existing solvers.
@@ 3.2 LLMEval
We observed that studies assessing language models’ factuality or evaluating whether the methods are effective to mitigate model hallucinations use different datasets and metrics. This makes it difficult to compare, in the same conditions, the factuality of different models as well as to compare the effectiveness of different factuality enhancement approaches. Moreover, a lot of prior work applied datasets such as MMLU (Hendrycks et al., 2021), StrategyQA (Geva et al., 2021) and HotpotQA (Yang and et al., 2018) to evaluate model’s factuality. These datasets tend to focus on evaluating the general performance, rather than factuality. To this end, we first collect a dataset FactQA by gathering a large number of factual questions that probe diverse factual errors and span across a spectrum of domains, to fairly evaluate LLMs’ factuality under the same criteria.
@@
We collect a set of questions by gathering questions from seven existing corpora that is collected deliberately to assess LLM’s factuality, including Snowball (Zhang et al., 2023a), SelfAware (Yin et al., 2023), FreshQA (Vu et al., 2023), FacTool (Chern et al., 2023), FELM-WK (Chen et al., 2023), Factcheck-GPT (Wang et al., 2023) and FactScore-Bio, a total of 6,480 examples shown in Table 1, referring to FactQA (see dataset details in Appendix A.1).
Each example includes the following fields: question, domain, topic, ability to test, task and source, where the domain and the topic are identified using GPT-4 based on the (question, reference response), and others are either inherited or summarized from the original datasets. [^1] We used GPT-4 response as a reference response for a question as it is more likely to provide a relevant and correct answer, assisting the identification of domains and topics.
Factual Error Type
@@
To concretely analyze models’ vulnerability, we tag three labels for each question from the perspective of the knowledge domain, the topic, and the potential error type if a LLM generates a factually incorrect response. Domains include a general domain, law, biomedical, clinical, science and so on. Given a domain, we further define fine-grained topics for the question. Three common error types are included.
@@
Type1: Knowledge error is the most common error, occurring when the model produces hallucinated or inaccurate information due to lacking relevant knowledge. That is, LLMs lack relevant knowledge or internalize false knowledge in the pre-training stage or in the problematic alignment process. However, LLMs do not know what they do not know, sometimes overestimate their capacities and confidently output unknown information, leading to false responses. Mitigating such errors require: (a) learning and correcting parametric knowledge through the curation of corpora used in pre-training, supervised fine-tuning (SFT) and alignment, (b) augmenting by external knowledge in inference, (c) calibrating models to be aware of unknowns, and (d) configuring the decoding strategies (sample/beam-search, temperature), balancing diversity and accuracy Zhang et al. (2023b); Wang et al. (2024).
@@
Type2: Over-commitment error occurs when the model fails to recognize the falsehoods (or jokes) inherent in the prompt or previously-generated context, and provides an inaccurate or inappropriate response. The left-to-right generation strategy used by LLMs poses potential risks that LLMs sometimes over-commit to the false premise in the context, even when they recognize they are incorrect (Zhang et al., 2023b). To address this issue, engineering better prompts is helpful, such as explicitly instructing models to first detect false premises in the prompt (Vu et al., 2023) and asking the same question in a different way (Is 10733 a prime number? $\rightarrow$ What are the factors of 10733? Let’s think step-by-step.)
@@
Type3: Disability error happens when the model is unable to search up-to-date information to correctly answer questions whose answers change over time, e.g., What is today’s gas price in New York (fast-changing), Who is the current spouse of Elon Musk (slow-changing). Retrieving external knowledge and augmenting it in the context would help.
@@
Note that we do not consider reasoning errors that arise when a claim employs flawed reasoning or faulty logic, and irrelevant error that denotes that the content is unrelated to the prompt (Chen et al., 2023). The former highlights LLM’s reasoning ability, which is more reflected in math and reasoning tasks, and the latter has more to do with response’s helpfulness or human preference. They are important in LLM evaluation, and may implicitly influence factuality, but we will first focus on explicit causes, leaving implicit claims for future work.
Evaluation Measurement
For questions that can be answered as Yes/No or have a short gold answer, we perform exact matching between the model responses and the gold standard answer to judge whether the response is factually correct or not, and then to calculate accuracy, such as for Snowball and SelfAware. There are no short gold answers in FreshQA, and thus we use the FreshEval proposed in Vu et al. (2023) to evaluate the correctness of model’s responses, in which few-shot in-context learning based on GPT-4 is used. We use a strict evaluation criterion, considering an answer to be correct only if all the claims in the response are factually true and also up-to-date.
For open-domain questions from the other four datasets with free-form and long responses, in which there are no gold standard answers, we use automatic fact-checking systems augmented with retrieved world-knowledge evidence to judge the correctness at the claim-level as well as at the document level.
@@
Table 2: The number of true, false, and unknown claims for FacTool-QA, FELM-WK and Factcheck-Bench, the number of responses for HaluEval (no claim-level labels). Unknown refers to no-enough-evidence claims and opinions.

### v2

# OpenFactCheck: Building, Benchmarking Customized Fact-Checking Systems and Evaluating the Factuality of Claims and LLMs

@@ Abstract
The increased use of large language models (LLMs) across a variety of real-world applications calls for mechanisms to verify the factual accuracy of their outputs. Difficulties lie in assessing the factuality of free-form responses in open domains. Also, different papers use disparate evaluation benchmarks and measurements, which renders them hard to compare and hampers future progress. To mitigate these issues, we propose OpenFactCheck, a unified framework for building customized automatic fact-checking systems, benchmarking their accuracy, evaluating factuality of LLMs, and verifying claims in a document. OpenFactCheck consists of three modules: (i) CustChecker allows users to easily customize an automatic fact-checker and verify the factual correctness of documents and claims, (ii) LLMEval, a unified evaluation framework assesses LLM’s factuality ability from various perspectives fairly, and (iii) CheckerEval is an extensible solution for gauging the reliability of automatic fact-checkers’ verification results using human-annotated datasets. Data and code are publicly available at https://github.com/yuxiaw/openfactcheck.
@@ 1 Introduction
Large language models (LLMs) have demonstrated impressive capabilities in generating naturally-sounding answers over a broad range of human inquiries. However, GPT-4 (OpenAI, 2023) and other text generation models still produce content that deviates from real-world facts (Bang et al., 2023; Wang et al., 2024b). This degrades the system performance and undermines its reliability, representing a significant bottleneck in the deployment (Chuang et al., 2024; Geng et al., 2024). We wonder about the percentage of factually incorrect content generated by LLMs and seek to identify the types of topics and scenarios in which LLMs are most prone to producing false claims.
@@
Many studies have explored evaluating and improving the factuality of LLMs Lee et al. (2022); Chuang et al. (2024); Shi et al. (2024); Chen et al. (2023). Two challenges have been identified for evaluation: (1) it is difficult to assess the factuality of open-domain free-form responses, and (2) different papers use different evaluation datasets and measurements, rendering them hard to compare and hampering future progress Wang et al. (2024b). To mitigate these issues, we introduce a fact-checking framework OpenFactCheck.
@@
Figure 1: Overview of the OpenFactCheck framework with three modules. Green CustChecker: a customized fact-checker to identify factual errors given the outputs of LLMs. Orange LLMEval: a unified LLM factuality evaluator to assess the LLM factual ability from different aspects, and then to produce a report highlighting the domains and types of factual errors the model frequently makes, along with the improvement advice. Purple CheckerEval: a fact-checker evaluator and leaderboard to encourage the development of advanced checkers in terms of performance, time and costs.
@@
It includes three modules as shown in Figure 1. CustChecker allows users to customize an automatic fact-checking system and to verify free-form documents to alleviate the first problem. LLMEval is a unified LLM factuality evaluation module. It applies seven factuality-specific benchmarks to assess the LLM factuality ability from different aspects and then produces a report to illustrate the weakness and offer improvement advice. This tackles the second challenge. We further incorporate CheckerEval that assesses the verification accuracy of automatic fact-checking systems, equipped with a leaderboard in terms of accuracy, latency, and costs, aiming to encourage the development of advanced automatic fact-checking systems.
@@
The three modules are designed to mutually reinforce one another. The human-annotated verification results derived from LLMEval can be used as the benchmark for evaluating the accuracy of automated fact-checkers. Simultaneously, the most effective checker identified in CheckerEval can be deployed for automated fact-checking tasks. Each fact-checker in CheckerEval can be an implementation in CustChecker. Complex user inquiries may be considered as potential candidates included into the factuality assessment dataset utilized in LLMEval.
@@
Users can tailor their checkers according to their specific needs, such as domain specialization, cost-effectiveness, or rapid processing, and identify factual errors for both human-written text (a claim or document) and the outputs of LLMs. LLM researchers and practitioners can directly input their LLM responses to the LLMEval, where responses are answers based on our question set. Subsequently, we conduct evaluations to assess the model’s factuality and to generate a report analyzing the model performance from multiple aspects, including snowballing hallucination, self-awareness of unknown knowledge, answering open-domain and latest-information questions. Similarly, developers who seek to evaluate and compare the efficacy of their fact-checking systems to other ones fairly can upload their checker’s verification outcomes to CheckerEval. Then, our system will show the ranking information in the leaderboard after evaluating under the same measurements.
To sum up, this work investigates three questions:
  • how to effectively identify factual errors in a document (e.g., an LLM response);
@@
We initiate an open-source project and develop a preliminary implementation of the three modules, which is anticipated to serve as a stepping stone to facilitate future endeavors in this domain. We encourage extensive implementation of unique, effective, and robust claim processors, retrievers and verifiers within fact-checking pipelines, collections of challenging questions where LLMs are prone to making factual errors, and human-annotated verification examples. We believe that this will help to promote future research on LLM factuality.
@@ 2 Background
In this section, we first analyze existing automatic fact-checking systems, followed by a review of previous approaches to evaluating the factuality of LLMs and the accuracy of fact-checking systems.
@@ 2.1 Fact-checking Systems
Fact-checking is the task of assessing whether claims made in writing are false or true. Many recent papers have described automatic fact-checking systems used to evaluate the factuality of LLM responses, such as RARR  Gao et al. (2023), FactScore Min et al. (2023), FacTool Chern et al. (2023), CoVe Dhuliawala et al. (2023), Factcheck-GPT Wang et al. (2024a) and FIRE Xie et al. (2024). Each checker has unique characteristics designed for specific domains or scenarios. RARR verifies a document as a whole and can generate an attribution report to explain factual errors. FactScore retrieves evidence from an offline Wikipedia dump mainly for the biography. FacTool is friendly to users with low latency, CoVe completely depends on the capability of LLMs. Factcheck-GPT has a fine-grained pipeline to localize intermediate errors, and FIRE is time- and cost-efficient. Unlike the above work, our aim is to enable the easy customization of a fact-checker according to users’ requirements and application scenarios, e.g., offline settings with a limited budget, by simply clicking dropdowns to choose offline retrievers and verifiers supported by small models without calling APIs.
@@
Despite different designs and implementations of various checkers, they generally consist of three modules: (1) claim processor, which extracts context-independent atomic claims from a document, (2) evidence retriever, which searches related passages from the Internet or database, and then ranks them by relevance, and (3) verifier, which determines the claim/document factuality based on the collected evidence (Guo et al., 2022; Li et al., 2024; Wang et al., 2024b). To this end, we first unify different fact-checking systems into a unified pipeline with the three modules. Then, we can select a developed module from various implementations or developers can implement one by themselves. In addition, the framework supports easy migration of existing fact-checking systems to our pipeline. However, the verification results of automatic fact-checkers are not necessarily accurate. How to evaluate and improve the accuracy of automated fact-checkers is critical, since the accuracy serves as a confidence and reliability signal for the verification results.
2.2 Evaluation of Fact-Checking Systems
@@
How accurate are current fact-checking systems? Can they effectively serve as proxies for evaluating the factual accuracy of language models? Existing automatic fact-checking studies often first collect a set of human-annotated (LLM response, extracted claims, factuality of the claims) in Figure 3, and then quantify the effectiveness of their systems by comparing the final verification results (i.e., whether a claim or a document is factually true or false) to human-annotated labels Min et al. (2023); Chern et al. (2023); Dhuliawala et al. (2023). Recent work on long-form factuality in LLMs also demonstrates a statistically significant correlation between the outputs by automatic fact-checkers and labels by human annotators (Wei et al., 2024). Thus, we merge four human-annotated LLM factuality datasets including FacTool-QA, FELM-WK, Factcheck-Bench, and HaluEval, and then compare them to the results of automatic fact-checkers to assess the performance of fact-checking systems.
@@ 2.3 LLM Factuality Evaluation
There are subtle differences between evaluating LLM’s general performance and factuality. Question answering (QA) datasets over various domains are always used for general performance evaluation (Hendrycks et al., 2021). The focus is on judging whether the question is answered correctly. If the model’s response contains the correct answer, it counts; otherwise, it is void even though the response presents all facts. While research on factuality primarily focuses on whether the response aligns with world knowledge, even if some statements were irrelevant to the question.
Therefore, instead of using datasets for general performance assessment, we selected seven datasets that are specifically collected for factuality evaluation. They were selected to cover as diverse potential factual errors as possible, including vulnerabilities of snowballing hallucinations, awareness of self-uncertainty, robustness to false-premise questions, fresh questions with answers changing fast over time, and free-form responses spanning distinct domains, topics, and tasks.
Summary
Observations above motivate us to integrate these three components into one framework to facilitate (i) users to flexibly configurate an automatic fact-checking system to verify the factuality of claims and documents, (ii) LLM developers to evaluate LLM’s factuality under the same measurement scale, and (iii) researchers to assess the fact-checkers reliability under fine-grained annotated benchmarks.
3 OpenFactCheck
OpenFactCheck’s design emphasizes two principles: (i) customizability and extensibility for both users and developers, and (ii) compatibility with existing methods and datasets. It consists of three modules: CustChecker, LLMEval, and CheckerEval. Below, we present the detailed design and implementation of each component.
@@ 3.1 CustChecker
CustChecker allows users to customize a fact-checking system by selecting a claim processor, a retriever, and a verifier. Current version supports the following fact-checking systems: RARR, FacTool and Factcheck-GPT (Gao et al., 2023; Chern et al., 2023; Wang et al., 2024a). Users can input either human-written text or outputs of LLMs, and then the fact-checker defined above will process and detect factual errors, showing the verification results including evidence, judgment, and explanations (see Figure 8).
@@
Table 1: FactQA: factual vulnerability, domain, potential error type and size across seven component datasets.
@@
We unify diverse fact-checking systems as a procedure with three steps, abstracted into three classes: a claim_processor, a retriever, and a verifier. The instances of the three classes are sequentially linked into a pipeline, solving the following tasks: (i) decomposing a document into atomic claims, (ii) collecting relevant evidence passages given a claim, and (iii) making a true/false judgment given both the claim and the evidence as input. We refer to them as task solvers. The implementation of a task solver can be flexible, just ensuring that the input and the output are aligned with the abstract class definitions. For example, evidence can be retrieved by calling SerperAPI or by searching Wikipedia using BM25, but we must return a list of relevant passages given an input claim.
Moreover, task solvers in our pipeline are not hardcoded, but can be configured through a yaml configuration file. Thus, users can combine task-solver implementations from different systems (e.g., using Factcheck-GPT’s claim processor, RARR’s retriever, and FacTool’s verifier) and start the verification from any step. For example, users can start from the step of retrieval when the input does not need decomposition.
This functionality is achieved by a message-passing mechanism, where a success flag is used to indicate whether the current task solver successfully executes and returns the expected output. The success flag passes through the configured solver order of the pipeline, guaranteeing that the output of the preceding solver fits the input for the current solver, otherwise error warning will be issued. Practically, the input and the output parameter names for the task solvers are defined in the configuration file. To link different solvers into a pipeline, one only needs to ensure that the current solver output name matches the input name of the succeeding solver. A dictionary format fact-checking-state is kept throughout the pipeline to store all information in the verification.
@@
Inspired by Fairseq, our framework is designed to be highly extendable by treating any third-party task solvers as plugins (Ott et al., 2019). As long as the developed task solvers adhere to our class interface definitions, they can be imported and used in our framework.
To sum, customizable and extendable nature of CustChecker allows general users to utilize it as an application with web-based user interfaces. Advanced developers have the flexibility to use it as a library, developing and integrating their solvers.
@@ 3.2 LLMEval
We observed that studies assessing language models’ factuality or evaluating whether the methods are effective to mitigate model hallucinations use different datasets and metrics. This makes it difficult to compare, in the same conditions, the factuality of different models as well as to compare the effectiveness of different factuality enhancement approaches. Moreover, a lot of prior work applied datasets such as MMLU (Hendrycks et al., 2021), StrategyQA (Geva et al., 2021) and HotpotQA (Yang and et al., 2018) to evaluate model’s factuality. These datasets tend to focus on assessing the general performance, rather than factuality. To this end, we first collect a dataset FactQA by gathering a large number of factual questions that probe diverse factual errors and span across a spectrum of domains, to fairly evaluate LLMs’ factuality under the same criteria.
@@
We collect a set of questions by gathering questions from seven existing corpora that is collected deliberately to assess LLM’s factuality, including Snowball (Zhang et al., 2024), SelfAware (Yin et al., 2023), FreshQA (Vu et al., 2024), FacTool-QA (Chern et al., 2023), FELM-WK (Chen et al., 2023), Factcheck-Bench (Wang et al., 2024a) and FactScore-Bio, a total of 6,480 examples shown in Table 1, referring to FactQA (see dataset details in Appendix A.1).
@@
To concretely analyze models’ vulnerability, we identify three labels for each question from the perspective of the knowledge domain, the topic, and the potential error type if a LLM generates a factually incorrect response. So each example includes the following fields: question, domain, topic, ability to test, task and source. Domain and topic are identified using GPT-4 based on the (question, reference response). [^1] We used GPT-4 response as a reference response for a question as it is more likely to provide a relevant and correct answer, assisting the identification of domains and topics. Domains involve general, legal, biomedical, clinical, scientific and so on. Given a domain, we further fine-grained topics. Three common error types are presented.
@@
Type1: Knowledge error is the most common error, occurring when the model produces hallucinated or inaccurate information due to lacking relevant knowledge or internalizing false knowledge in the pre-training stage or in the problematic alignment process. However, LLMs do not know what they do not know, sometimes overestimate their capacities and confidently output unknown information, leading to false responses. Mitigating such errors require: (a) learning and correcting parametric knowledge through the curation of corpora used in pre-training, supervised fine-tuning (SFT) and alignment, (b) augmenting by external knowledge in inference, (c) calibrating models to be aware of unknowns, and (d) configuring the decoding strategies (sample/beam-search, temperature), balancing diversity and accuracy Zhang et al. (2023).
@@
Type2: Over-commitment error occurs when the model fails to recognize the falsehoods (or jokes) inherent in the prompt or previously-generated context, and provides an inaccurate or inappropriate response. The left-to-right generation strategy used by LLMs poses potential risks that LLMs sometimes over-commit to the false premise in the context, even when they recognize they are incorrect (Zhang et al., 2023). To address this issue, engineering better prompts is helpful, such as explicitly instructing models to first detect false premises in the prompt (Vu et al., 2024) and asking the same question in a different way (Is 10733 a prime number? $\rightarrow$ What are the factors of 10733? Let’s think step-by-step.)
@@
Type3: Disability error happens when the model is unable to search up-to-date information to correctly answer questions whose answers change over time, e.g., What is today’s gas price in New York (fast-changing). Retrieving external knowledge and augmenting it in the context would help.
@@
Note that we do not consider reasoning errors that arise when a claim employs flawed reasoning or faulty logic, and irrelevant error concerning that the content is unrelated to the question (Chen et al., 2023). The former highlights LLM’s reasoning ability, which is more reflected in math and reasoning tasks, and the latter has more to do with response’s helpfulness or human preference. They are important in LLM evaluation, and may implicitly influence factuality, but we will first focus on explicit causes, leaving the implicit for future work.
@@
Table 2: The number of true, false claims and unknown (no-enough-evidence or opinions) for FacTool-QA, FELM-WK and Factcheck-Bench, the number of responses for HaluEval (no claim-level labels).
Evaluation Measurement
For questions that can be answered by Yes/No or have a short gold answer, we perform exact matching between the model responses and the gold standard answer to judge whether the response is factually correct or not, and then to calculate accuracy, such as for Snowball and SelfAware. For FreshQA, we use the FreshEval proposed in Vu et al. (2024) to evaluate the correctness of model’s responses, in which few-shot in-context learning based on GPT-4 is applied. We use the strict evaluation criterion which considers an answer to be correct only if all the claims in the response are factually true and also up-to-date.
For open-domain questions from the other four datasets with free-form and long responses, there are no gold standard answers. We use automatic fact-checking systems augmented with retrieved world-knowledge evidence to judge the correctness at the claim-level as well as the document level.